### v1

# Unknotting Lagrangian $\mathrm{S}^{1}\times\mathrm{S}^{n-1}$ in $\mathbb{R}^{2n}$

Abstract.
Lagrangian embeddings $\mathrm{S}^{1}\times\mathrm{S}^{n-1}\hookrightarrow\mathbb{R}^{2n}$ are classified up to smooth isotopy for all $n\geq 3$.
This work was partially supported by the SFB TRR 191 Symplectic Structures in Geometry, Algebra and Dynamics, funded by the DFG (Projektnummer 281071066 – TRR 191).
1. Introduction and overview
Two embeddings of a manifold $N$ in a manifold $M$ are isotopic if they can be connected by a family of embeddings. Diffeomorphisms of $N$ act on the set of isotopy classes of embeddings $N\hookrightarrow M$ by precomposition. The quotient by this action is the set of isotopy classes of (unparametrised) submanifolds in $M$ diffeomorphic to $N$.
The purpose of this paper is to describe the isotopy classes represented by Lagrangian embeddings of $\mathrm{S}^{1}\times\mathrm{S}^{n-1}$ in $(\mathbb{R}^{2n},\omega_{\mathrm{st}})$. The case $n=1$ is elementary, so we assume $n\geq 2$. $\mathbb{R}^{2n}$ is identified with $\mathbb{C}^{n}$ so that the symplectic form is given by the imaginary part of the standard Hermitian metric. An embedding $j:N\to\mathbb{C}^{n}$ is Lagrangian if multiplication by $i=\sqrt{-1}$ maps the tangent space of $j(N)$ at each point onto its Euclidean orthogonal complement. Lagrangian embeddings are obviously totally real, that is, have no tangent complex lines.
An explicit example of a Lagrangian embedding $\mathrm{S}^{1}\times\mathrm{S}^{n-1}\hookrightarrow\mathbb{C}^{n}$ for every $n$ is given by
$\mathrm{S}^{1}\times\mathrm{S}^{n-1}\ni(\theta,x)\longmapsto(1+e^{i\theta}/2)\bigl{(}x_{1},...,x_{n}\bigr{)}\in\mathbb{C}^{n},$ (1.1)
where we are viewing $\mathrm{S}^{1}$ as $\mathbb{R}/2\pi\mathbb{Z}$ and $\mathrm{S}^{n-1}$ as the unit sphere in $\mathbb{R}^{n}$.
Our main Theorem 1.1 shows that for $n\neq 2,4$ every Lagrangian embedding $\mathrm{S}^{1}\times\mathrm{S}^{n-1}\hookrightarrow\mathbb{C}^{n}$ is isotopic to (1.1) if $n$ is even and either to (1.1) or to its composition with the reflection $(\theta,x)\mapsto(-\theta,x)$ if $n$ is odd. The isotopy is not through Lagrangian embeddings in general. In fact, there exist Lagrangian embeddings $\mathrm{S}^{1}\times\mathrm{S}^{n-1}\hookrightarrow\mathbb{C}^{n}$ for all $n>2$ that are not isotopic through Lagrangian embeddings to any reparametrisation of (1.1), see [18, Theorem 5] and [8, Corollary 1.6].
Theorem 1.1.
If $n\geq 6$ is even, then all Lagrangian embeddings of $\mathrm{S}^{1}\times\mathrm{S}^{n-1}$ in $\mathbb{C}^{n}$ are smoothly isotopic. If $n\geq 3$ is odd, then Lagrangian embeddings of $\mathrm{S}^{1}\times\mathrm{S}^{n-1}$ in $\mathbb{C}^{n}$ fall into two distinct smooth isotopy classes related by precomposition with a reflection of $\mathrm{S}^{1}$.
The result for $n=8$ was obtained by Borrelli [3]; our argument is a generalisation of his approach based on [17, §6]. Dimitroglou Rizell and Evans [5, Corollary E] proved the result for Lagrangian embeddings with minimal Maslov number $n$ for even $n\geq 6$ and for monotone Lagrangian embeddings for odd $n\geq 5$. However, embeddings (1.1) have minimal Maslov number $2$ for all $n$ and non-monotone Lagrangian embeddings $\mathrm{S}^{1}\times\mathrm{S}^{2k}\to\mathbb{C}^{2k+1}$ exist for all $k\geq 1$ by [8, Corollary 1.6].
The two cases $n=2$ and $n=4$ excluded from the theorem are of a rather different nature:
[…]
Corollary 1.2.
All Lagrangian submanifolds in $\mathbb{C}^{n}$ diffeomorphic to $\mathrm{S}^{1}\times\mathrm{S}^{n-1}$ are smoothly isotopic if $n\neq 4$.
The extent to which Theorem 1.1 and Corollary 1.2 restrict the differential topology of Lagrangian embeddings and submanifolds for $n\geq 3$ can be seen from the following brief summary of embedding theory:
[…]
Theorem 1.1 is a typical ‘symplectic rigidity’ result. It becomes completely false if Lagrangian embeddings are replaced by ‘soft’ totally real embeddings in $\mathbb{C}^{n}$. Namely, every embedding $\mathrm{S}^{1}\times\mathrm{S}^{n-1}\hookrightarrow\mathbb{C}^{n}$ is smoothly isotopic to a totally real embedding. To see this, recall that any two embeddings of an orientable $n$-manifold in $\mathbb{C}^{n}$ are regularly homotopic through immersions. (This follows from the general theory in [14].) Therefore the tangent map of any embedding of $\mathrm{S}^{1}\times\mathrm{S}^{n-1}$ is homotopic to the totally real tangent map of an (existing) Lagrangian embedding. An isotopy to a totally real embedding is then given by the $h$-principle [11, §19.3].
The proof of Theorem 1.1 follows the scheme from [3] and consists of two parts discussed in Sections 2 and 3 respectively:
(1) For a totally real embedding $j:\mathrm{S}^{1}\times\mathrm{S}^{n-1}\to\mathbb{C}^{n}$ and a nowhere vanishing vector field $v$ tangent to the $\mathrm{S}^{1}$-fibres, we may consider the pushoff of $j$ in the direction of $i\cdot dj(v)$. If the embedding is Lagrangian, this pushoff is nullhomologous in $\mathbb{C}^{n}\setminus j(\mathrm{S}^{1}\times\mathrm{S}^{n-1})$ or, equivalently, is unlinked with $(n-1)$-cycles in $\mathrm{S}^{1}\times\mathrm{S}^{n-1}$ up to replacing $v$ with $-v$ for odd $n$. The proof of this assertion uses a generalisation of the Luttinger surgery [15] to arbitrary dimensions in the same way as the original surgery was used in [10]. Such a surgery produces (for $n\geq 3$) an exact symplectic manifold that is symplectomorphic to $(\mathbb{R}^{2n},\omega_{\mathrm{st}})$ outside of a compact subset and must therefore have the homology of $\mathbb{R}^{2n}$ by a basic application of pseudoholomorphic curves [16]. Comparing this with the homology calculations in the explicit model from [17, §6] proves the vanishing of the relevant linking numbers.
(2) Embeddings of an $n$-manifold $N$ in $\mathbb{R}^{2n}$ for $n\geq 4$ are governed by an $h$-principle established by Haefliger and Hirsch [13]. Isotopy classes of such embeddings are in one-to-one correspondence with homotopy classes of certain equivariant maps from the frame bundle of $N^{\circ}$ to the Stiefel manifold $\mathcal{V}_{2n,n+1}$, where $N^{\circ}$ is the complement to a small ball in $N$. The additional $(n+1)$-st component of the map is given by a normal vector field such that the pushoff of $N$ is unlinked with $N^{\circ}$. In particular, for a Lagrangian embedding of $N=\mathrm{S}^{1}\times\mathrm{S}^{n-1}$, this normal field can be taken to be $i\cdot dj(\pm v)$ by part (1). (The sign only makes a difference for odd $n$ and can be adjusted by a reflection of $\mathrm{S}^{1}$.) If we now recall that our manifold is parallelisable and our embedding is totally real, the relevant frame map reduces to a map from $N^{\circ}$ itself of the form $x\mapsto\Phi(x)\gamma(x)$, where $\Phi(x)\in\mathrm{GL}(n,\mathbb{C})$ and $\gamma$ is a fixed map to the Stiefel manifold. Homotopy calculations (conveniently done in [5]) show that for $n\neq 2,4$ the map $\Phi$ has no effect on the homotopy class, and the result for $n>4$ follows. In the remaining case $n=3$, a version of the same argument in the setting of [19] determines the value of the Whitney invariant $W$, which turns out to be ±$\pm 1$, so that there is only one possible isotopy class up to a reflection of $\mathrm{S}^{1}$.
2. Generalised Luttinger surgery, symplectic rigidity, and the linking class
Represent $\Sigma_{n}:=\mathrm{S}^{1}\times\mathrm{S}^{n-1}$, $n\geq 2$, as the quotient of $\mathbb{R}^{n}\setminus\{0\}$ by the $\mathbb{Z}$-action generated by the transformation
$x\longmapsto 2x.$ (2.1)
The cotangent bundle $T^{*}\Sigma_{n}$ with its canonical symplectic structure is the quotient of $T^{*}(\mathbb{R}^{n}\setminus\{0\})\cong(\mathbb{R}^{n}_{x}\setminus\{0\})\times\mathbb{R}^{n}_{y}$ with the symplectic form $\sum dx_{k}\wedge dy_{k}$ by the symplectic $\mathbb{Z}$-action generated by
$(x,y)\longmapsto(2x,\frac{1}{2}y)$ (2.2)
in the ‘complex’ coordinates $x=q$ and $y=-p$ on $T^{*}\mathbb{R}^{n}$.
The Riemannian metric
$g=\frac{1}{\|x\|^{2}}{\sum dx_{k}^{2}}$
on $\mathbb{R}^{n}\setminus\{0\}$ is invariant with respect to (2.1) and therefore induces a metric on $\Sigma_{n}$. ($\Sigma_{n}$ with this metric is the Riemannian product of the standard unit sphere in $\mathbb{R}^{n}$ and the circle of length $\log 2$.) Note further that the bundle $S_{r}^{*}(\mathbb{R}^{n}\setminus\{0\})\subset T^{*}(\mathbb{R}^{n}\setminus\{0\})$ of cotangent spheres of radius $r>0$ with respect to $g$ is the hypersurface
$\bigl{\{}\|x\|^{2}\|y\|^{2}=r^{2}\bigr{\}}\subset(\mathbb{R}^{n}_{x}\setminus\{0\})\times(\mathbb{R}^{n}_{y}\setminus\{0\}).$
On $T^{*}(\mathbb{R}^{n}\setminus\{0\})$ with the zero section removed, consider the map
$(x,y)\longmapsto(-y,x).$ (2.3)
Obviously, this map is a symplectomorphism preserving $S_{r}^{*}(\mathbb{R}^{n}\setminus\{0\})$ for every $r>0$. Furthermore, it maps the orbits of the action (2.2) into orbits. Hence, it defines a symplectomorphism $f$ of $T^{*}\Sigma_{n}$ with the zero section removed which maps $S_{r}^{*}\Sigma_{n}$ into itself for every $r>0$.
Similarly, the $g$-isometry
$x\longmapsto\frac{x}{\|x\|^{2}}$ (2.4)
acts on $\Sigma_{n}=\mathrm{S}^{1}\times\mathrm{S}^{n-1}$ by a reflection of $\mathrm{S}^{1}$ and its co-differential defines a symplectomorphism $\tau$ of the entire $T^{*}\Sigma_{n}$ preserving $S_{r}^{*}\Sigma_{n}$.
Let $j:\Sigma_{n}\to\mathbb{C}^{n}$ be a Lagrangian embedding and identify $\Sigma_{n}$ with the zero section of $T^{*}\Sigma_{n}$. By the Weinstein tubular neighbourhood theorem, $j$ extends to a symplectomorphism from the disc bundle $D_{r}^{*}\Sigma_{n}$ with $r>0$ small enough onto a closed tubular neighbourhood $\overline{U}$ of $L=j(\Sigma_{n})\subset\mathbb{C}^{n}$. For any diffeomorphism $h$ of $S_{r}^{*}\Sigma_{n}\cong\partial U$ from the subgroup generated by $f$ and $\tau$, consider the manifold
$X=X(j,h):=\overline{U}\cup_{h}\mathbb{C}^{n}\setminus U$
with the symplectic form $\omega_{X}$ equal to the standard symplectic form of $\mathbb{C}^{n}$ on $\overline{U}$ and $\mathbb{C}^{n}\setminus U$. This form is well-defined because $h$ extends to a symplectomorphism of $\overline{U}\setminus L$.
Lemma 2.1.
The symplectic manifold $(X,\omega_{X})$ is exact for $n\geq 3$.
Proof.
Consider the following exact sequence in compactly supported cohomology
$\dots\to\mathrm{H}^{2n-2}_{c}(X\setminus L;\mathbb{R})\to\mathrm{H}^{2n-2}_{c}(X;\mathbb{R})\to\mathrm{H}^{2n-2}_{c}(L;\mathbb{R})\to\dots$
The last term vanishes for $n\geq 3$ and hence the first map is surjective. $X\setminus L$ is symplectomorphic to $\mathbb{C}^{n}\setminus L$, which is exact, so $[\omega_{X}]$ pairs trivially with $\mathrm{H}^{2n-2}_{c}(X\setminus L;\mathbb{R})$. Thus, $[\omega_{X}]$ pairs trivially with $\mathrm{H}^{2n-2}_{c}(X;\mathbb{R})$ and therefore $\omega_{X}$ is exact by Poincaré duality in cohomology. ∎
Remark 2.2.
For $n\geq 4$, the sequence
$0\cong\mathrm{H}^{2n-3}_{c}(L;\mathbb{R})\to\mathrm{H}^{2n-2}_{c}(\mathbb{C}^{n}\setminus L;\mathbb{R})\to\mathrm{H}^{2n-2}_{c}(\mathbb{C}^{n};\mathbb{R})\cong 0$
shows that $\mathrm{H}^{2n-2}_{c}(X\setminus L;\mathbb{R})=\mathrm{H}^{2n-2}_{c}(\mathbb{C}^{n}\setminus L;\mathbb{R})=0$ and then the sequence in the proof of the lemma implies $\mathrm{H}^{2n-2}_{c}(X;\mathbb{R})=0$. Hence, $\mathrm{H}^{2}(X;\mathbb{R})=0$ by duality, and this conclusion remains valid for any orientable Dehn surgery on $L$ as in [17, Lemma 9]. Thus, the argument using the symplectic structure on $X$ is only needed for $n=3$.
Proposition 2.3.
$X$ is diffeomorphic to $\mathbb{C}^{n}$.
Proof.
For $n\geq 3$, this is a consequence of Lemma 2.1 and the following classical rigidity result of Eliashberg, Floer, and McDuff [9, 16]: A symplectically aspherical (e.g. exact) symplectic manifold symplectomorphic to $(\mathbb{C}^{n},\omega_{\mathrm{st}})$ outside of a compact subset is diffeomorphic to $\mathbb{C}^{n}$. For $n=2$, Luttinger [15, p. 222] observed that $X$ is even symplectomorphic to $(\mathbb{C}^{2},\omega_{\mathrm{st}})$. ∎
Remark 2.4.
We shall actually use the weaker assertion that $X$ has the homology of $\mathbb{C}^{n}$ for $n\geq 3$. This follows directly from [16, §3.8] and Lemma 2.1.
Take the orientation on $\mathbb{R}^{n}$ by the order of the coordinates and orient $T^{*}\mathbb{R}^{n}$ and $\mathbb{C}^{n}$ as $\mathbb{R}^{n}\oplus i\mathbb{R}^{n}$. (This orientation differs from the standard complex/symplectic one by the sign $(-1)^{\frac{1}{2}n(n-1)}$, which will make some of the following formulas nicer.)
In the $(n-1)$-dimensional homology group $\mathrm{H}_{n-1}(S_{r}^{*}\Sigma_{n};\mathbb{Z})$, consider the classes of the fibre
$\delta:=\bigl{[}\{\|y\|=r,x=(1,0,\dots,0)\}\bigr{]}$
and of the ‘horizontal’ sphere
$\gamma:=\bigl{[}\{\|x\|=1,y=(r,0,\dots,0)\}\bigr{]}$
oriented as the boundaries of the corresponding balls in $\mathbb{R}^{n}$.
Lemma 2.5.
Let $f$ and $\tau$ be the diffeomophisms of $S_{r}^{*}\Sigma_{n}$ defined above. Then
$\left\{\begin{array}[]{rcl}f_{*}\delta&=&(-1)^{n}\gamma\\ f_{*}\gamma&=&\delta\end{array}\right.\quad\text{ and }\quad\left\{\begin{array}[]{rcl}\tau_{*}\delta&=&-\delta\\ \tau_{*}\gamma&=&\gamma+(1+(-1)^{n})\delta\end{array}\right.$
and the same identities hold in homology with coefficients in $\mathbb{Z}/p\mathbb{Z}$.
Proof.
The first three formulas follow immediately from the definitions and the fact that the antipodal map reverses the orientation on $\mathrm{S}^{n-1}$ if and only if $n$ is odd. To get the formula for $\tau_{*}\gamma$, note that
$\begin{array}[]{rcl}\tau_{*}(\gamma+\delta)&=&\bigl{[}\tau\bigl{(}\{\|x\|=1,y=rx\}\bigr{)}\bigr{]}\\[2.0pt] &=&\bigl{[}\{\|x\|=1,y=-rx\}\bigr{]}\\[2.0pt] &=&\gamma+(-1)^{n}\delta\end{array}$
and combine this with $\tau_{*}\delta=-\delta$. ∎
Recall that if $N\subset\mathbb{C}^{n}$ is a connected oriented closed submanifold of dimension $k$, then the linking number
$\mathop{\mathrm{lk}}\nolimits(N,\cdot):\mathrm{H}_{2n-1-k}(\mathbb{C}^{n}\setminus N;\mathbb{Z})\to\mathbb{Z}$
is an isomorphism. Furthermore, its reduction $\mathrel{\mathrm{mod}}p$ defines an isomorphism $\mathrm{H}_{2n-1-k}(\mathbb{C}^{n}\setminus N;\mathbb{Z}/p\mathbb{Z})\overset{\cong}{\longrightarrow}\mathbb{Z}/p\mathbb{Z}$ for every $p$.
Proposition 2.6.
Let $j:\Sigma_{n}\overset{\cong}{\longrightarrow}L\subset\mathbb{C}^{n}$ be a Lagrangian embedding and let $\widetilde{\jmath}:S_{r}^{*}\Sigma_{n}\overset{\cong}{\longrightarrow}\partial U\subset\mathbb{C}^{n}$ be the induced embedding of the boundary of a Weinstein neighbourhood $U$. Then $\mathop{\mathrm{lk}}\nolimits(L,\widetilde{\jmath}_{*}\gamma)=\pm 1$ and moreover $\mathop{\mathrm{lk}}\nolimits(L,\widetilde{\jmath}_{*}\gamma)=-1$ if $n$ is even and the orientations are chosen as above.
Proof.
If $\mathop{\mathrm{lk}}\nolimits(L,\widetilde{\jmath}_{*}\gamma)=0\mathrel{\mathrm{mod}}p$, then $-(-1)^{n}\widetilde{\jmath}_{*}\gamma$ bounds an $n$-chain $\mathrel{\mathrm{mod}}p$ in $\mathbb{C}^{n}\setminus U$. This chain and the $n$-ball bounded by $\widetilde{\jmath}_{*}\delta$ in $\overline{U}$ form an $n$-cycle $\mathrel{\mathrm{mod}}p$ in $X=X(j,f)=\overline{U}\cup_{f}(\mathbb{C}^{n}\setminus U)$ because $f_{*}\delta=(-1)^{n}\gamma$ by Lemma 2.5. The intersection index of $L$ with this cycle in $X$ is equal to $1$ and hence $L$ represents a non-trivial class in $\mathrm{H}_{n}(X;\mathbb{Z}/p\mathbb{Z})$, which contradicts Proposition 2.3. So $\mathop{\mathrm{lk}}\nolimits(L,\widetilde{\jmath}_{*}\gamma)$ is non-zero mod every $p$, which means that it equals ±$\pm 1$.
If $n$ is even, we may also consider $X=X(j,\tau f)$. Then $\tau_{*}f_{*}\delta=\gamma+2\delta$ by Lemma 2.5 and, taking $p=3$ in the preceding argument, we obtain
$0\neq\mathop{\mathrm{lk}}\nolimits(L,\widetilde{\jmath}_{*}(\gamma+2\delta))\mathrel{\mathrm{mod}}3=\mathop{\mathrm{lk}}\nolimits(L,\widetilde{\jmath}_{*}\gamma)+2\mathrel{\mathrm{mod}}3,$
which excludes the value $1$. ∎
Let $j:N\to\mathbb{C}^{n}$ be a totally real (e.g. Lagrangian) embedding of a compact manifold and $\zeta\in\Gamma(TN)$ a nowhere vanishing vector field on $N$. The $\mathbb{C}$-normal pushoff of $j$ by $\zeta$ is the isotopy class in $\mathbb{C}^{n}\setminus j(N)$ of the embedding
$j^{\sharp}_{\zeta}(a)=j(a)+\varepsilon i\cdot dj_{a}(\zeta(a)),\quad a\in N,$ (2.5)
for a sufficiently small $\varepsilon>0$. If $N$ is closed and oriented, then the homology class
$\sigma(j,\zeta):=[j^{\sharp}_{\zeta}(N)]\in\mathrm{H}_{n}(\mathbb{C}^{n}\setminus j(N);\mathbb{Z})$ (2.6)
is called the linking class of $j$ and $\zeta$, cf. [3, §2].
If $\psi$ is a diffeomorphism of $N$, then
$(j\circ\psi)^{\sharp}_{\zeta}(a)=j(\psi(a))+\varepsilon i\cdot dj_{\psi(a)}(d\psi_{a}(\zeta(a)))=(j^{\sharp}_{\psi_{*}\zeta}\circ\psi)(a)$
and therefore
$\sigma(j\circ\psi,\zeta)=\pm\sigma(j,\psi_{*}\zeta),$ (2.7)
where the sign depends on whether $\psi$ preserves or reverses the orientation on $N$.
Theorem 2.7.
Let $j:\Sigma_{n}\to\mathbb{C}^{n}$ be a Lagrangian embedding and take the nowhere vanishing vector field $v=\sum x_{k}\frac{\partial}{\partial x_{k}}$ on $\Sigma_{n}=\mathbb{R}_{x}^{n}\setminus\{0\}/\mathbb{Z}$.
  • 1)
    If $n$ is even, then $\sigma(j,v)=0$.
  • 2)
    If $n$ is odd, then either $\sigma(j,v)=0$ or $\sigma(j,-v)=0$.
Remark 2.8.
This theorem extends to arbitrary $n\geq 2$ the results of Eliashberg–Polterovich [10] for $n=2$ and Borrelli [3] for $n=4,8$.
Proof.
Let $m$ denote the meridional $(n-1)$-sphere $\{\|x\|=1\}\subset\Sigma_{n}$. In our model for the Weinstein neighbourhood of $L=j(\Sigma_{n})$, the $\mathbb{C}$-normal pushoffs $j^{\sharp}_{\pm v}(m)$ of $m$ by ±$\pm v$ are represented by $\{\|x\|=1,y=\pm rx\}$, so
$\bigl{[}j^{\sharp}_{\pm v}(m)\bigr{]}=\widetilde{\jmath}_{*}(\gamma+(\pm 1)^{n}\delta)\in\mathrm{H}_{n-1}(\mathbb{C}^{n}\setminus L;\mathbb{Z}).$
Now we can apply Proposition 2.6:
  • If $n$ is even, then
    $\mathop{\mathrm{lk}}\nolimits\bigl{(}L,[j^{\sharp}_{-v}(m)]\bigr{)}=\mathop{\mathrm{lk}}\nolimits\bigl{(}L,\widetilde{\jmath}_{*}(\gamma+\delta)\bigr{)}=-1+1=0.$ (2.8)
  • If $n$ is odd, there are two cases. If $\mathop{\mathrm{lk}}\nolimits(L,\widetilde{\jmath}_{*}(\gamma))=1$, then
    $\mathop{\mathrm{lk}}\nolimits(L,[j^{\sharp}_{-v}(m)])=\mathop{\mathrm{lk}}\nolimits\bigl{(}L,\widetilde{\jmath}_{*}(\gamma-\delta)\bigr{)}=1-1=0.$ (2.9)
    If $\mathop{\mathrm{lk}}\nolimits(L,\widetilde{\jmath}_{*}(\gamma))=-1$, then
    $\mathop{\mathrm{lk}}\nolimits(L,[j^{\sharp}_{v}(m)])=\mathop{\mathrm{lk}}\nolimits\bigl{(}L,\widetilde{\jmath}_{*}(\gamma+\delta)\bigr{)}=-1+1=0.$ (2.10)
Note also that by (2.6), (2.5), and the (anti-)commutation rule for the linking number (see e.g. [7, p. 234]), we have
$\begin{array}[]{rcl}\mathop{\mathrm{lk}}\nolimits\bigl{(}j(m),\sigma(j,w)\bigr{)}&=&\mathop{\mathrm{lk}}\nolimits\bigl{(}j(m),[j^{\sharp}_{w}(\Sigma_{n})]\bigr{)}\\[2.0pt] &=&\mathop{\mathrm{lk}}\nolimits\bigl{(}j^{\sharp}_{-w}(m),[j(\Sigma_{n})]\bigr{)}\\[2.0pt] &=&\mathop{\mathrm{lk}}\nolimits\bigl{(}L,[j^{\sharp}_{-w}(m)]\bigr{)}\end{array}$ (2.11)
for any nowhere vanishing vector field $w$ on $\Sigma_{n}$.
Hence, $\sigma(j,v)$ is unlinked with $j(m)$ by (2.11) and (2.8) for even $n$. For odd $n$, either $\sigma(j,v)$ is unlinked with $j(m)$ by (2.11) and (2.9) or $\sigma(j,-v)$ is unlinked with $j(m)$ by (2.11) and (2.10).
If $n\geq 3$, the linking number with $j(m)$ defines an isomorphism $\mathrm{H}_{n}(\mathbb{C}^{n}\setminus j(\Sigma_{n});\mathbb{Z})\cong\mathbb{Z}$, which completes the proof.
If $n=2$, then $\mathrm{H}_{2}(\mathbb{C}^{2}\setminus j(\mathrm{S}^{1}\times\mathrm{S}^{1});\mathbb{Z})\cong\mathbb{Z}^{2}$ and we also have to consider the linking number with the longitudinal circle $\ell=\mathrm{S}^{1}\times\{\mathrm{pt}\}$. Let $\psi$ be the obvious diffeomorphism of $\mathrm{S}^{1}\times\mathrm{S}^{1}$ swapping $m$ and $\ell$. Applying the above argument to $j\circ\psi$, we obtain
[…]
However, the vector field $\psi_{*}v\perp v$ is homotopic to $v$ through nowhere vanishing vector fields on $\mathrm{S}^{1}\times\mathrm{S}^{1}$ and therefore $\sigma(j,\psi_{*}v)=\sigma(j,v)$ by (2.6) and (2.5). Hence, $\sigma(j,v)$ is unlinked with $j(\ell)$ too. ∎
Corollary 2.9.
If $\sigma(j,v)\neq 0$ for a Lagrangian embedding $j:\Sigma_{n}\to\mathbb{C}^{n}$, then $n$ is odd and $\sigma(j\circ\rho,v)=0$ for the diffeomorphism $\rho$ given by (2.4).
Proof.
$\rho_{*}v=-v$ and the claim follows from (2.7) and Theorem 2.7. ∎
3. Classification of embeddings
Let us start with the special case $n=3$. The isotopy classes of smooth embeddings $j:\Sigma_{3}\to\mathbb{R}^{6}\cong\mathbb{C}^{3}$ are indexed by the following two invariants [19, p. 648]. First, the Whitney invariant
$W(j)\in\mathrm{H}_{1}(\Sigma_{3};\mathbb{Z})\cong\mathbb{Z},$
which is essentially the difference class between the Haefliger–Hirsch maps (3.2) of $j$ and of the standard embedding in $\mathbb{R}^{2}\times\mathbb{R}^{3}\cong\mathbb{R}^{5}\subset\mathbb{R}^{6}$. Second, the Kreck invariant
$\eta(j)\in\mathbb{Z}/|W(j)|\mathbb{Z},$
where we have used the above identification $\mathrm{H}_{1}(\Sigma_{3};\mathbb{Z})\cong\mathbb{Z}$. The Kreck invariant generalises the Haefliger invariant [12] classifying smoothly knotted $3$-spheres in $\mathbb{R}^{6}$.
In particular and most importantly for us, the Kreck invariant of an embedding with $W(j)=\pm 1\in\mathbb{Z}$ takes values in $\{0\}$ and therefore all such embeddings with the same Whitney invariant are isotopic. (This case is singled out in Corollary (a) on p. 649 of [19].)
The Boéchat–Haefliger Invariant Lemma on p. 661 of [19] asserts that if $\xi$ is an unlinked section of the normal bundle of $j$, then
$2W(j)=\pm\mathop{\mathrm{PD}}\,e(\xi^{\perp}))\in\mathrm{H}_{1}(\Sigma_{3};\mathbb{Z}),$ (3.1)
that is, $2W(j)$ is Poincaré dual to the Euler class of the orthogonal complement of $\xi$ in the normal bundle of $j(\Sigma_{3})$. The sign $\pm$ in (3.1) is fixed by orientation conventions as stated right before the Difference Lemma on p. 658 of [19].
For a Lagrangian embedding $j$, an unlinked normal bundle section is given by $\xi:=i\cdot dj(v)$ after possibly precomposing $j$ with the diffeomorphism (2.4) by Corollary 2.9. Multiplication by $i$ defines an isomorphism $\xi^{\perp}\cong v^{\perp}$, where the latter complement is taken in $T\Sigma_{3}$. The orthogonal complement to $v$ is just the subbundle tangent to the $\mathrm{S}^{2}$-fibres in $\Sigma_{3}$, so its Euler class is clearly equal to twice a generator of $\mathrm{H}^{2}(\Sigma_{3};\mathbb{Z})$, whence
$W(j)=\pm 1\in\mathbb{Z}\cong\mathrm{H}_{1}(\Sigma_{3};\mathbb{Z})$
by formula (3.1), where the sign is again fixed by the orientation conventions and the identification $\mathrm{H}_{1}(\Sigma_{3};\mathbb{Z})\cong\mathbb{Z}$. Now the above remark about the Kreck invariant implies the following result for $n=3$.
Proposition 3.1.
All Lagrangian embeddings $\mathrm{S}^{1}\times\mathrm{S}^{2}\hookrightarrow\mathbb{C}^{3}$ are smoothly isotopic up to precomposition with (2.4).
Let now $n\geq 2$ be arbitrary. Trivialise the tangent bundle of $\Sigma_{n}$ using the dilation invariant vector fields
$\zeta_{k}=\|x\|\frac{\partial}{\partial x_{k}},\quad k=1,\ldots,n,$
on $\mathbb{R}^{n}\setminus\{0\}$ and note that
$v=\sum\frac{x_{k}}{\|x\|}\zeta_{k}.$
Fix also a complex linear trivialisation of $T\mathbb{C}^{n}$ so that $T_{a}\mathbb{C}^{n}\cong\mathbb{C}^{n}$ for all $a\in\mathbb{C}^{n}$.
Let $j:\Sigma_{n}\to\mathbb{C}^{n}$ be an embedding and let $\Sigma_{n}^{\circ}$ denote the complement of a small $n$-ball in $\Sigma_{n}$. There exists a unique up to homotopy normal vector field $\nu_{j}$ on $j(\Sigma_{n})$ such that it does not vanish on $\Sigma_{n}^{\circ}$ and the pushoff of $j(\Sigma)$ by $\nu_{j}$ is nullhomologous in $\mathbb{C}^{n}\setminus j(\Sigma_{n}^{\circ})$, see [13, p. 134]. (This vector field is called the associated vector field of $j$ in [3, §5.1] and the Haefliger–Hirsch vector field in [5, §6.2]. It is also the unlinked section of the normal bundle from [19] appearing in (3.1).)
Specialising [13, Theorem (2.2)] and the remarks right after it to our situation, we see that for $n\geq 4$ the isotopy classes of embeddings $j$ are in one-to-one correspondence with the homotopy classes of their Haefliger–Hirsch maps
$\Sigma_{n}^{\circ}\ni x\longmapsto\bigl{(}dj_{x}(\zeta_{1}(x)),\dots,dj_{x}(\zeta_{n}(x)),\nu_{j}(j(x)\bigr{)}\in\mathcal{V}_{2n,n+1}$ (3.2)
to the Stiefel manifold of real $(n+1)$-frames in $\mathbb{R}^{2n}\cong\mathbb{C}^{n}$, cf. [3, §5.2]. (It is convenient to use arbitrary frames here but it is well-known that $\mathcal{V}_{m,\ell}$ retracts onto the Stiefel manifold of orthonormal frames [20, §2].)
Since $\mathcal{V}_{2n,n+1}$ is $(n-2)$-connected [20, Satz 8] and $\Sigma_{n}^{\circ}$ retracts onto the wedge sum $\mathrm{S}^{1}\vee\mathrm{S}^{n-1}$, it follows that for $n\geq 3$ the homotopy class of (3.2) is determined by the image of the meridional $(n-1)$-sphere $m\subset\Sigma_{n}^{\circ}$ in $\pi_{n-1}(\mathcal{V}_{2n,n+1})$. The latter group is isomorphic to $\mathbb{Z}$ if $n$ is odd and to $\mathbb{Z}/2\mathbb{Z}$ if $n$ is even [20, Satz 9], which completes the classification of smooth embeddings of $\Sigma_{n}$ in $\mathbb{C}^{n}$ for $n\geq 4$.
Assume now that $j$ is a Lagrangian embedding. By Corollary 2.9, up to precomposing $j$ with the reflection (2.4) for odd $n$, we may choose $\nu_{j}=i\cdot dj(v)$ so that the map (3.2) takes the form
$x\longmapsto\left(dj_{x}(\zeta_{1}(x)),\dots,dj_{x}(\zeta_{n}(x)),i\sum\limits_{k=1}^{n}\frac{x_{k}}{\|x\|}dj_{x}(\zeta_{k}(x))\right)$
Furthermore, since a Lagrangian embedding is totally real, the vectors $dj_{x}(\zeta_{k}(x))$ form a complex frame of $\mathbb{C}^{n}$ for every $x\in\Sigma_{n}^{\circ}$. Therefore there is a smooth map
$\Phi:\Sigma_{n}^{\circ}\longrightarrow\mathrm{GL}(n,\mathbb{C})$
such that
$dj_{x}(\zeta_{k}(x))=\Phi(x)e_{k},$
where $\{e_{k}\}_{k=1,\dots,n}$ is our fixed complex frame of $\mathbb{C}^{n}$. Hence, the map (3.2) can be written as
$x\longmapsto\left(\Phi(x)e_{1},\dots,\Phi(x)e_{n},\Phi(x)\left(\sum\limits_{k=1}^{n}\frac{x_{k}}{\|x\|}ie_{k}\right)\right).$ (3.3)
The action of $\mathrm{GL}(n,\mathbb{C})\subset\mathrm{GL}_{+}(2n,\mathbb{R})$ on $\mathcal{V}_{2n,n+1}$ induces a trivial map on the relevant homotopy group $\pi_{n-1}$ for all $n\neq 2,4$ by the proof of [5, Lemma 6.14]. Specifically,
  • (1)
    if $n$ is odd, then $\pi_{n-1}(\mathrm{GL}(n,\mathbb{C}))=\pi_{n-1}(\mathrm{U}(n))=0$;
  • (2)
    if $n\neq 2,4,8$ is even, then the map
    $\pi_{n-1}(\mathrm{GL}_{+}(2n,\mathbb{R}))\longrightarrow\pi_{n-1}(\mathcal{V}_{2n,n+1})=\mathbb{Z}/2\mathbb{Z}$
    induced by the action is trivial;
  • (3)
    if $n=8$, then the map in (2) is surjective but the map
    $\mathbb{Z}=\pi_{7}(\mathrm{GL}(8,\mathbb{C}))\longrightarrow\pi_{7}(\mathrm{GL}_{+}(16,\mathbb{R}))=\mathbb{Z}$
    is the multiplication by $2$, as observed already in [3, §5.3].
Thus, for $n\neq 2,4$, the Haefliger–Hirsch map (3.3) of any Lagrangian embedding (precomposed with (2.4) if $n$ is odd and $\sigma(j,v)\neq 0$) is homotopic to the standard map
$x\longmapsto\left(e_{1},\dots,e_{n},\sum\limits_{k=1}^{n}\frac{x_{k}}{\|x\|}ie_{k}\right),$ (3.4)
which proves the following result for all $n\geq 5$.
Proposition 3.2.
All Lagrangian embeddings $\mathrm{S}^{1}\times\mathrm{S}^{n-1}\hookrightarrow\mathbb{C}^{n}$, $n\geq 5$, are smoothly isotopic up to precomposition with (2.4) if $n$ is odd.
Remark 3.3.
If $n=4$, the following argument from [3, p. 232] shows that both isotopy classes of embeddings $j:\Sigma_{4}\to\mathbb{C}^{4}$ have Lagrangian representatives. Note first that (for any $n\geq 3$) the action of the map $\Phi$ in (3.3) on $\pi_{n-1}(\Sigma_{n}^{\circ})$ can be chosen at will by [1, §6.2] and the $h$-principle for Lagrangian immersions of $\mathrm{S}^{n-1}$. However, the map
$\mathbb{Z}=\pi_{3}(\mathrm{GL}(4,\mathbb{C}))\longrightarrow\pi_{3}(\mathcal{V}_{8,5})=\mathbb{Z}/2\mathbb{Z}$
is surjective and so each homotopy class of maps $\Sigma_{4}^{\circ}\to\mathcal{V}_{8,5}$ can be represented by the Haefliger–Hirsch map of a Lagrangian embedding.
It remains to show that precomposition with the involution $\rho$ of $\Sigma_{n}$ defined by (2.4) does indeed change the smooth isotopy class of a Lagrangian embedding for odd $n$. Swapping $j$ and $j\circ\rho$ if necessary, we may assume that $\sigma(j,v)=0$ by Corollary 2.9 so that $\sigma(j\circ\rho,-v)=0$ by (2.7). Hence, the argument leading to Proposition 3.2 shows that the Haefliger–Hirsch maps of $j$ and $j\circ\rho$ are homotopic to
$x\longmapsto\left(e_{1},\dots,e_{n},\pm\sum\limits_{k=1}^{n}\frac{x_{k}}{\|x\|}ie_{k}\right)$ (3.5)
with the $+$ sign for $j$ and the -- sign for $j\circ\rho$. The restrictions of the maps (3.5) to the meridional sphere $m=\{\|x\|=1\}$ parametrise the (homotopy) fibre $\mathrm{S}^{n-1}\sim\mathcal{V}_{n,1}$ of the ‘forgetful’ fibre bundle
$\mathcal{V}_{2n,n+1}\longrightarrow\mathcal{V}_{2n,n}$
and differ by the antipodal map of $\mathrm{S}^{n-1}$. For odd $n$, the antipodal map is orientation reversing and hence acts non-trivially on $\pi_{n-1}(\mathrm{S}^{n-1})=\mathbb{Z}$. Furthermore, $\pi_{n-1}(\mathcal{V}_{2n,n})=0$ by [20, Satz 8] and $\pi_{n}(\mathcal{V}_{2n,n})=\mathbb{Z}/2\mathbb{Z}$ for odd $n>1$ by [20, Satz 9], so the homotopy exact sequence
$\pi_{n}(\mathcal{V}_{2n,n})\to\pi_{n-1}(\mathrm{S}^{n-1})\to\pi_{n-1}(\mathcal{V}_{2n,n+1})\to\pi_{n-1}(\mathcal{V}_{2n,n})$
shows that the embedding of the fibre is an isomorphism on $\pi_{n-1}$. Hence, the two maps in (3.5) are not homotopic, which proves the next statement.
Proposition 3.4.
A Lagrangian embedding $\mathrm{S}^{1}\times\mathrm{S}^{n-1}\hookrightarrow\mathbb{C}^{n}$ with odd $n\geq 3$ is not smoothly isotopic to its composition with (2.4).
Remark 3.5.
For $n=3$ one can alternatively (and in fact equivalently) use formula (3.1) and the specific choice of the unlinked section $\xi$ in the proof of Proposition 3.1 to show that $W(j\circ\rho)=-W(j)$ for a Lagrangian embedding $j$.
Theorem 1.1 follows by combining Propositions 3.1, 3.2, and 3.4.