### v3

# Unknotting Lagrangian $\mathrm{S}^{1}\times\mathrm{S}^{n-1}$ in $\mathbb{R}^{2n}$

Abstract.
Lagrangian embeddings $\mathrm{S}^{1}\times\mathrm{S}^{n-1}\hookrightarrow\mathbb{R}^{2n}$ are classified up to smooth isotopy for all $n\geq 3$.
This work was partially supported by the SFB TRR 191 Symplectic Structures in Geometry, Algebra and Dynamics, funded by the DFG (Projektnummer 281071066 – TRR 191).
1. Introduction and overview
Two embeddings of a manifold $N$ in a manifold $M$ are isotopic if they can be connected by a family of embeddings. Diffeomorphisms of $N$ act on the set of isotopy classes of embeddings $N\hookrightarrow M$ by precomposition. The quotient by this action is the set of isotopy classes of (unparametrised) submanifolds in $M$ diffeomorphic to $N$.
The purpose of this paper is to describe the isotopy classes represented by Lagrangian embeddings of $\mathrm{S}^{1}\times\mathrm{S}^{n-1}$ in $(\mathbb{R}^{2n},\omega_{\mathrm{st}})$. We assume $n\geq 2$ to exclude the elementary special case $n=1$. $\mathbb{R}^{2n}$ is identified with $\mathbb{C}^{n}$ so that the symplectic form is the imaginary part of the standard Hermitian metric. An embedding $j:N\to\mathbb{C}^{n}$ is Lagrangian if multiplication by $i=\sqrt{-1}$ maps the tangent space of $j(N)$ at each point onto its Euclidean orthogonal complement. Lagrangian embeddings are obviously totally real, that is, have no tangent complex lines.
An explicit example of a Lagrangian embedding $\mathrm{S}^{1}\times\mathrm{S}^{n-1}\hookrightarrow\mathbb{C}^{n}$ for every $n$ is given by
$\mathrm{S}^{1}\times\mathrm{S}^{n-1}\ni\;(\theta,x_{1},\dots,x_{n})\longmapsto\bigl{(}1+\frac{1}{2}e^{i\theta}\bigr{)}(x_{1},\dots,x_{n})\;\in\mathbb{C}^{n},$ (1.1)
where we are viewing $\mathrm{S}^{1}$ as $\mathbb{R}/2\pi\mathbb{Z}$ and $\mathrm{S}^{n-1}$ as the unit sphere in $\mathbb{R}^{n}$.
Our main Theorem 1.1 shows that for $n\neq 2,4$ every Lagrangian embedding $\mathrm{S}^{1}\times\mathrm{S}^{n-1}\hookrightarrow\mathbb{C}^{n}$ is isotopic to (1.1) if $n$ is even and either to (1.1) or to its composition with the reflection $(\theta,x)\mapsto(-\theta,x)$ if $n$ is odd. The isotopy is not through Lagrangian embeddings in general. In fact, there exist Lagrangian embeddings $\mathrm{S}^{1}\times\mathrm{S}^{n-1}\hookrightarrow\mathbb{C}^{n}$ for all $n>2$ that are not isotopic through Lagrangian embeddings to any reparametrisation of (1.1), see [18, §4], [9, Corollary 1.6], and [19, §4].
Theorem 1.1.
If $n\geq 6$ is even, then all Lagrangian embeddings of $\mathrm{S}^{1}\times\mathrm{S}^{n-1}$ in $\mathbb{C}^{n}$ are smoothly isotopic. If $n\geq 3$ is odd, then Lagrangian embeddings of $\mathrm{S}^{1}\times\mathrm{S}^{n-1}$ in $\mathbb{C}^{n}$ fall into two distinct smooth isotopy classes related by precomposition with a reflection of $\mathrm{S}^{1}$.
The result for $n=8$ was obtained by Borrelli [3]; our argument is a generalisation of his approach based on [17, §6]. Dimitroglou Rizell and Evans [6, Corollary E] proved the result for Lagrangian embeddings with minimal Maslov number $n$ for even $n\geq 6$ and for monotone Lagrangian embeddings for odd $n\geq 5$. However, embeddings (1.1) have minimal Maslov number $2$ for all $n$ and non-monotone Lagrangian embeddings $\mathrm{S}^{1}\times\mathrm{S}^{2k}\to\mathbb{C}^{2k+1}$ exist for all $k\geq 1$ by [9, Corollary 1.6].
The two cases $n=2$ and $n=4$ excluded from the theorem are of a rather different nature:
[…]
Corollary 1.2.
If $n\neq 4$, then all Lagrangian submanifolds in $\mathbb{C}^{n}$ diffeomorphic to $\mathrm{S}^{1}\times\mathrm{S}^{n-1}$ are smoothly isotopic to the image of (1.1).
The extent to which Theorem 1.1 and Corollary 1.2 restrict the differential topology of Lagrangian embeddings and submanifolds for $n\geq 3$ can be seen from the following brief summary of embedding theory:
[…]
Theorem 1.1 is a typical ‘symplectic rigidity’ result. It becomes completely false if Lagrangian embeddings are replaced by ‘soft’ totally real embeddings in $\mathbb{C}^{n}$. Namely, every embedding $\mathrm{S}^{1}\times\mathrm{S}^{n-1}\hookrightarrow\mathbb{C}^{n}$ is smoothly isotopic to a totally real embedding. To see this, recall that any two embeddings of an orientable $n$-manifold in $\mathbb{C}^{n}$ are regularly homotopic through immersions. (This follows from the general theory in [14].) Therefore the tangent map of any embedding of $\mathrm{S}^{1}\times\mathrm{S}^{n-1}$ is homotopic to the totally real tangent map of an (existing) Lagrangian embedding. An isotopy to a totally real embedding is then given by the $h$-principle [4, §27.4].
The proof of Theorem 1.1 follows the scheme from [3] and consists of two parts discussed in Section 2 and Section 3, respectively:
(1) For a totally real embedding $j:\mathrm{S}^{1}\times\mathrm{S}^{n-1}\to\mathbb{C}^{n}$ and a nowhere vanishing vector field $v$ tangent to the $\mathrm{S}^{1}$-fibres, we may consider the pushoff of $j$ in the direction of $i\cdot dj(v)$. If the embedding is Lagrangian, this pushoff is nullhomologous in $\mathbb{C}^{n}\setminus j(\mathrm{S}^{1}\times\mathrm{S}^{n-1})$ or, equivalently, is unlinked with $(n-1)$-cycles in $\mathrm{S}^{1}\times\mathrm{S}^{n-1}$ up to replacing $v$ with $-v$ for odd $n$. The proof of this assertion uses a generalisation of the Luttinger surgery [15] to arbitrary dimensions in the same way as the original surgery was used in [11]. Such a surgery produces (for $n\geq 3$) an exact symplectic manifold that is symplectomorphic to $(\mathbb{R}^{2n},\omega_{\mathrm{st}})$ outside of a compact subset and that must therefore have the homology of $\mathbb{R}^{2n}$ by a basic application of pseudoholomorphic curves [16]. Comparing this with the homology calculations in the explicit model from [17, §6] proves the vanishing of the relevant linking numbers.
(2) Embeddings of a connected orientable $n$-manifold $N$ in $\mathbb{R}^{2n}$ for $n\geq 4$ are governed by an $h$-principle due to Haefliger and Hirsch [13]. Isotopy classes of such embeddings are in one-to-one correspondence with homotopy classes of equivariant maps from the frame bundle of $N^{\circ}$ to the Stiefel manifold $\mathcal{V}_{2n,n+1}$, where $N^{\circ}$ is the complement to a small ball in $N$. The $(n+1)$-st component of the map to $\mathcal{V}_{2n,n+1}$ is given by a normal vector field on $N$ such that the pushoff of $N$ is unlinked with $N^{\circ}$. Thus, for a Lagrangian embedding of $N=\mathrm{S}^{1}\times\mathrm{S}^{n-1}$, this normal field can be taken to be $i\cdot dj(\pm v)$ by part (1). (The sign only makes a difference for odd $n$ and can be adjusted by a reflection of $\mathrm{S}^{1}$.) Recalling now that our manifold is parallelisable and our embedding is totally real, we can reduce the frame map to a map from $N^{\circ}$ itself of the form $x\mapsto\Phi(x)\beta(x)$, where $\Phi(x)\in\mathrm{GL}(n,\mathbb{C})$ and $\beta$ is a fixed map to the Stiefel manifold. Homotopy calculations (conveniently done in [6]) show that for $n\neq 2,4$ the map $\Phi$ has no effect on the homotopy class, and the result for $n>4$ follows. In the remaining case $n=3$, a version of the same argument in the setting of [20] shows that the Whitney invariant $W$ equals ±$\pm 1$ so that $\mathbb{Z}/|W|\mathbb{Z}=\{0\}$ and there is only one possible isotopy class up to a reflection of $\mathrm{S}^{1}$.
2. Generalised Luttinger surgery, symplectic rigidity, and the linking class
Represent $\Sigma_{n}:=\mathrm{S}^{1}\times\mathrm{S}^{n-1}$, $n\geq 2$, as the quotient of $\mathbb{R}^{n}\setminus\{0\}$ by the $\mathbb{Z}$-action generated by the transformation
$x\longmapsto 2x.$ (2.1)
The cotangent bundle $T^{*}\Sigma_{n}$ with its canonical symplectic structure is the quotient of $T^{*}(\mathbb{R}^{n}\setminus\{0\})\cong(\mathbb{R}^{n}_{x}\setminus\{0\})\times\mathbb{R}^{n}_{y}$ with the symplectic form $\sum dx_{k}\wedge dy_{k}$ by the symplectic $\mathbb{Z}$-action generated by
$(x,y)\longmapsto(2x,\frac{1}{2}y)$ (2.2)
in the ‘complex’ coordinates $x=q$ and $y=-p$ on $T^{*}\mathbb{R}^{n}$.
The Riemannian metric
$g=\frac{1}{\|x\|^{2}}{\sum dx_{k}^{2}}$
on $\mathbb{R}^{n}\setminus\{0\}$ is invariant with respect to (2.1) and therefore induces a metric on $\Sigma_{n}$. ($\Sigma_{n}$ with this metric is the Riemannian product of the standard unit sphere in $\mathbb{R}^{n}$ and the circle of length $\log 2$.) Note further that the bundle $S_{r}^{*}(\mathbb{R}^{n}\setminus\{0\})\subset T^{*}(\mathbb{R}^{n}\setminus\{0\})$ of cotangent spheres of radius $r>0$ with respect to $g$ is the hypersurface
$\bigl{\{}\|x\|^{2}\|y\|^{2}=r^{2}\bigr{\}}\subset(\mathbb{R}^{n}_{x}\setminus\{0\})\times(\mathbb{R}^{n}_{y}\setminus\{0\}).$
On $T^{*}(\mathbb{R}^{n}\setminus\{0\})$ with the zero section removed, consider the map
$(x,y)\longmapsto(-y,x).$ (2.3)
Obviously, this map is a symplectomorphism preserving $S_{r}^{*}(\mathbb{R}^{n}\setminus\{0\})$ for every $r>0$. Furthermore, it maps the orbits of the action (2.2) into orbits. Hence, it defines a symplectomorphism $f$ of $T^{*}\Sigma_{n}$ with the zero section removed which maps $S_{r}^{*}\Sigma_{n}$ into itself for every $r>0$.
Similarly, the $g$-isometry
$x\longmapsto\frac{x}{\|x\|^{2}}$ (2.4)
acts on $\Sigma_{n}=\mathrm{S}^{1}\times\mathrm{S}^{n-1}$ by a reflection of $\mathrm{S}^{1}$ and its co-differential defines a symplectomorphism $\tau$ of the entire $T^{*}\Sigma_{n}$ preserving $S_{r}^{*}\Sigma_{n}$.
Let $j:\Sigma_{n}\to\mathbb{C}^{n}$ be a Lagrangian embedding and identify $\Sigma_{n}$ with the zero section of $T^{*}\Sigma_{n}$. By the Weinstein tubular neighbourhood theorem, $j$ extends to a symplectomorphism from the disc bundle $D_{r}^{*}\Sigma_{n}$ with $r>0$ small enough onto a closed tubular neighbourhood $\overline{U}$ of $L=j(\Sigma_{n})\subset\mathbb{C}^{n}$. For any diffeomorphism $h$ of $S_{r}^{*}\Sigma_{n}\cong\partial U$ from the subgroup generated by $f$ and $\tau$, consider the manifold
$X=X(j,h):=\overline{U}\cup_{h}\mathbb{C}^{n}\setminus U$
with the symplectic form $\omega_{X}$ equal to the standard symplectic form of $\mathbb{C}^{n}$ on $\overline{U}$ and $\mathbb{C}^{n}\setminus U$. This form is well-defined because $h$ extends to a symplectomorphism of $\overline{U}\setminus L$.
Lemma 2.1.
The symplectic manifold $(X,\omega_{X})$ is exact for $n\geq 3$.
Proof.
Consider the following exact sequence in compactly supported cohomology
$\dots\to\mathrm{H}^{2n-2}_{c}(X\setminus L;\mathbb{R})\to\mathrm{H}^{2n-2}_{c}(X;\mathbb{R})\to\mathrm{H}^{2n-2}_{c}(L;\mathbb{R})\to\dots$
The last term vanishes for $n\geq 3$ and hence the first map is surjective. $X\setminus L$ is symplectomorphic to $\mathbb{C}^{n}\setminus L$, which is exact, so $[\omega_{X}]$ pairs trivially with $\mathrm{H}^{2n-2}_{c}(X\setminus L;\mathbb{R})$. Thus, $[\omega_{X}]$ pairs trivially with $\mathrm{H}^{2n-2}_{c}(X;\mathbb{R})$ and therefore $\omega_{X}$ is exact by Poincaré duality in cohomology. ∎
Remark 2.2.
For $n\geq 4$, the sequence
$0\cong\mathrm{H}^{2n-3}_{c}(L;\mathbb{R})\to\mathrm{H}^{2n-2}_{c}(\mathbb{C}^{n}\setminus L;\mathbb{R})\to\mathrm{H}^{2n-2}_{c}(\mathbb{C}^{n};\mathbb{R})\cong 0$
shows that $\mathrm{H}^{2n-2}_{c}(X\setminus L;\mathbb{R})=\mathrm{H}^{2n-2}_{c}(\mathbb{C}^{n}\setminus L;\mathbb{R})=0$ and then the sequence in the proof of the lemma implies $\mathrm{H}^{2n-2}_{c}(X;\mathbb{R})=0$. Hence, $\mathrm{H}^{2}(X;\mathbb{R})=0$ by duality, and this conclusion remains valid for any orientable Dehn surgery on $L$ as in [17, Lemma 9]. Thus, the argument using the symplectic structure on $X$ is only needed for $n=3$.
Proposition 2.3.
$X$ is diffeomorphic to $\mathbb{C}^{n}$.
Proof.
For $n\geq 3$, this is a consequence of Lemma 2.1 and the following classical rigidity result of Eliashberg, Floer, and McDuff [10, 16]: A symplectically aspherical (e.g. exact) symplectic manifold symplectomorphic to $(\mathbb{C}^{n},\omega_{\mathrm{st}})$ outside of a compact subset is diffeomorphic to $\mathbb{C}^{n}$. For $n=2$, Luttinger [15, p. 222] observed that $X$ is even symplectomorphic to $(\mathbb{C}^{2},\omega_{\mathrm{st}})$. ∎
Remark 2.4.
We shall actually use the weaker assertion that $X$ has the homology of $\mathbb{C}^{n}$ for $n\geq 3$. This follows directly from [16, §3.8] and Lemma 2.1.
Take the orientation on $\mathbb{R}^{n}$ by the order of the coordinates and orient $T^{*}\mathbb{R}^{n}$ and $\mathbb{C}^{n}$ as $\mathbb{R}^{n}\oplus i\mathbb{R}^{n}$. (This orientation differs from the standard complex/symplectic one by the sign $(-1)^{n(n-1)/2}$, which will make some of the following formulas nicer.)
In the $(n-1)$-dimensional homology group $\mathrm{H}_{n-1}(S_{r}^{*}\Sigma_{n};\mathbb{Z})$, consider the classes of the fibre
$\delta:=\bigl{[}\{\|y\|=r,x=(1,0,\dots,0)\}\bigr{]}$
and of the ‘horizontal’ sphere
$\gamma:=\bigl{[}\{\|x\|=1,y=(r,0,\dots,0)\}\bigr{]}$
oriented as the boundaries of the corresponding balls in $\mathbb{R}^{n}$.
Lemma 2.5.
Let $f$ and $\tau$ be the diffeomorphisms of $S_{r}^{*}\Sigma_{n}$ defined above. Then
$\left\{\begin{array}[]{rcl}f_{*}\delta&=&(-1)^{n}\gamma\\ f_{*}\gamma&=&\delta\end{array}\right.\quad\text{ and }\quad\left\{\begin{array}[]{rcl}\tau_{*}\delta&=&-\delta\\ \tau_{*}\gamma&=&\gamma+(1+(-1)^{n})\delta\end{array}\right.$
and the same identities hold in homology with coefficients in $\mathbb{Z}/p\mathbb{Z}$.
Proof.
The first three formulas follow immediately from the definitions and the fact that the antipodal map reverses the orientation on $\mathrm{S}^{n-1}$ if and only if $n$ is odd. To get the formula for $\tau_{*}\gamma$, note that
$\begin{array}[]{rcl}\tau_{*}(\gamma+\delta)&=&\bigl{[}\tau\bigl{(}\{\|x\|=1,y=rx\}\bigr{)}\bigr{]}\\[2.0pt] &=&\bigl{[}\{\|x\|=1,y=-rx\}\bigr{]}\\[2.0pt] &=&\gamma+(-1)^{n}\delta\end{array}$
and combine this with $\tau_{*}\delta=-\delta$. ∎
Recall that if $N\subset\mathbb{C}^{n}$ is a connected oriented closed submanifold of dimension $k$, then the linking number
$\mathop{\mathrm{lk}}\nolimits(N,\cdot):\mathrm{H}_{2n-1-k}(\mathbb{C}^{n}\setminus N;\mathbb{Z})\to\mathbb{Z}$
is an isomorphism. Furthermore, its reduction $\mathrel{\mathrm{mod}}p$ defines an isomorphism $\mathrm{H}_{2n-1-k}(\mathbb{C}^{n}\setminus N;\mathbb{Z}/p\mathbb{Z})\overset{\cong}{\longrightarrow}\mathbb{Z}/p\mathbb{Z}$ for every $p$.
Proposition 2.6.
Let $j:\Sigma_{n}\overset{\cong}{\longrightarrow}L\subset\mathbb{C}^{n}$ be a Lagrangian embedding and let $\widetilde{\jmath}:S_{r}^{*}\Sigma_{n}\overset{\cong}{\longrightarrow}\partial U\subset\mathbb{C}^{n}$ be the induced embedding of the boundary of a Weinstein neighbourhood $U$. Then $\mathop{\mathrm{lk}}\nolimits(L,\widetilde{\jmath}_{*}\gamma)=\pm 1$ and moreover $\mathop{\mathrm{lk}}\nolimits(L,\widetilde{\jmath}_{*}\gamma)=-1$ if $n$ is even and the orientations are chosen as above.
Proof.
If $\mathop{\mathrm{lk}}\nolimits(L,\widetilde{\jmath}_{*}\gamma)=0\mathrel{\mathrm{mod}}p$, then $-(-1)^{n}\widetilde{\jmath}_{*}\gamma$ bounds an $n$-chain $\mathrel{\mathrm{mod}}p$ in $\mathbb{C}^{n}\setminus U$. This chain and the $n$-ball bounded by $\widetilde{\jmath}_{*}\delta$ in $\overline{U}$ form an $n$-cycle $\mathrel{\mathrm{mod}}p$ in $X=X(j,f)=\overline{U}\cup_{f}(\mathbb{C}^{n}\setminus U)$ because $f_{*}\delta=(-1)^{n}\gamma$ by Lemma 2.5. The intersection index of $L$ with this cycle in $X$ is equal to $1$ and hence $L$ represents a non-trivial class in $\mathrm{H}_{n}(X;\mathbb{Z}/p\mathbb{Z})$, which contradicts Proposition 2.3. So $\mathop{\mathrm{lk}}\nolimits(L,\widetilde{\jmath}_{*}\gamma)$ is non-zero mod every $p$, which means that it equals ±$\pm 1$.
If $n$ is even, we may also consider $X=X(j,\tau f)$. Then $\tau_{*}f_{*}\delta=\gamma+2\delta$ by Lemma 2.5 and, taking $p=3$ in the preceding argument, we obtain
$0\neq\mathop{\mathrm{lk}}\nolimits(L,\widetilde{\jmath}_{*}(\gamma+2\delta))\mathrel{\mathrm{mod}}3=\mathop{\mathrm{lk}}\nolimits(L,\widetilde{\jmath}_{*}\gamma)+2\mathrel{\mathrm{mod}}3,$
which excludes the value $1$. ∎
Let $j:N\to\mathbb{C}^{n}$ be a totally real (e.g. Lagrangian) embedding of a compact manifold and $\zeta\in\Gamma(TN)$ a nowhere vanishing vector field on $N$. The $\mathbb{C}$-normal pushoff of $j$ by $\zeta$ is the isotopy class in $\mathbb{C}^{n}\setminus j(N)$ of the embedding
$j^{\sharp}_{\zeta}(a)=j(a)+\varepsilon i\cdot dj_{a}(\zeta(a)),\quad a\in N,$ (2.5)
for a sufficiently small $\varepsilon>0$. If $N$ is closed and oriented, then the homology class
$\sigma(j,\zeta):=[j^{\sharp}_{\zeta}(N)]\in\mathrm{H}_{n}(\mathbb{C}^{n}\setminus j(N);\mathbb{Z})$ (2.6)
is called the linking class of $j$ and $\zeta$, cf. [3, §2].
If $\psi$ is a diffeomorphism of $N$, then
$(j\circ\psi)^{\sharp}_{\zeta}(a)=j(\psi(a))+\varepsilon i\cdot dj_{\psi(a)}(d\psi_{a}(\zeta(a)))=(j^{\sharp}_{\psi_{*}\zeta}\circ\psi)(a)$
and therefore
$\sigma(j\circ\psi,\zeta)=\pm\sigma(j,\psi_{*}\zeta),$ (2.7)
where the sign depends on whether $\psi$ preserves or reverses the orientation on $N$.
Theorem 2.7.
Let $j:\Sigma_{n}\to\mathbb{C}^{n}$ be a Lagrangian embedding and take the nowhere vanishing vector field $v=\sum x_{k}\frac{\partial}{\partial x_{k}}$ on $\Sigma_{n}=\left(\mathbb{R}_{x}^{n}\setminus\{0\}\right)/\mathbb{Z}$.
  • 1)
    If $n$ is even, then $\sigma(j,v)=0$.
  • 2)
    If $n$ is odd, then either $\sigma(j,v)=0$ or $\sigma(j,-v)=0$.
Remark 2.8.
This theorem extends to arbitrary $n\geq 2$ the results of Eliashberg–Polterovich [11] for $n=2$ and Borrelli [3] for $n=4,8$.
Proof.
Let $m$ denote the meridional $(n-1)$-sphere $\{\|x\|=1\}\subset\Sigma_{n}$. In our model for the Weinstein neighbourhood of $L=j(\Sigma_{n})$, the $\mathbb{C}$-normal pushoffs $j^{\sharp}_{\pm v}(m)$ of $m$ by ±$\pm v$ are represented by $\{\|x\|=1,y=\pm rx\}$, so
$\bigl{[}j^{\sharp}_{\pm v}(m)\bigr{]}=\widetilde{\jmath}_{*}(\gamma+(\pm 1)^{n}\delta)\in\mathrm{H}_{n-1}(\mathbb{C}^{n}\setminus L;\mathbb{Z}).$
Now we can apply Proposition 2.6:
  • If $n$ is even, then
    $\mathop{\mathrm{lk}}\nolimits\bigl{(}L,[j^{\sharp}_{-v}(m)]\bigr{)}=\mathop{\mathrm{lk}}\nolimits\bigl{(}L,\widetilde{\jmath}_{*}(\gamma+\delta)\bigr{)}=-1+1=0.$ (2.8)
  • If $n$ is odd, there are two cases. If $\mathop{\mathrm{lk}}\nolimits(L,\widetilde{\jmath}_{*}(\gamma))=1$, then
    $\mathop{\mathrm{lk}}\nolimits(L,[j^{\sharp}_{-v}(m)])=\mathop{\mathrm{lk}}\nolimits\bigl{(}L,\widetilde{\jmath}_{*}(\gamma-\delta)\bigr{)}=1-1=0.$ (2.9)
    If $\mathop{\mathrm{lk}}\nolimits(L,\widetilde{\jmath}_{*}(\gamma))=-1$, then
    $\mathop{\mathrm{lk}}\nolimits(L,[j^{\sharp}_{v}(m)])=\mathop{\mathrm{lk}}\nolimits\bigl{(}L,\widetilde{\jmath}_{*}(\gamma+\delta)\bigr{)}=-1+1=0.$ (2.10)
Note also that by (2.6), (2.5), and the (anti-)commutation rule for the linking number (see e.g. [5, p. 234]), we have
$\begin{array}[]{rcl}\mathop{\mathrm{lk}}\nolimits\bigl{(}j(m),\sigma(j,w)\bigr{)}&=&\mathop{\mathrm{lk}}\nolimits\bigl{(}j(m),[j^{\sharp}_{w}(\Sigma_{n})]\bigr{)}\\[2.0pt] &=&\mathop{\mathrm{lk}}\nolimits\bigl{(}j^{\sharp}_{-w}(m),[j(\Sigma_{n})]\bigr{)}\\[2.0pt] &=&\mathop{\mathrm{lk}}\nolimits\bigl{(}L,[j^{\sharp}_{-w}(m)]\bigr{)}\end{array}$ (2.11)
for any nowhere vanishing vector field $w$ on $\Sigma_{n}$.
Hence, $\sigma(j,v)$ is unlinked with $j(m)$ by (2.11) and (2.8) for even $n$. For odd $n$, either $\sigma(j,v)$ is unlinked with $j(m)$ by (2.11) and (2.9) or $\sigma(j,-v)$ is unlinked with $j(m)$ by (2.11) and (2.10).
If $n\geq 3$, the linking number with $j(m)$ defines an isomorphism $\mathrm{H}_{n}(\mathbb{C}^{n}\setminus j(\Sigma_{n});\mathbb{Z})\cong\mathbb{Z}$, which completes the proof.
If $n=2$, then $\mathrm{H}_{2}(\mathbb{C}^{2}\setminus j(\mathrm{S}^{1}\times\mathrm{S}^{1});\mathbb{Z})\cong\mathbb{Z}^{2}$ and we also have to consider the linking number with the longitudinal circle $\ell=\mathrm{S}^{1}\times\{\mathrm{pt}\}$. Let $\psi$ be the obvious diffeomorphism of $\mathrm{S}^{1}\times\mathrm{S}^{1}$ swapping $m$ and $\ell$. Applying the above argument to $j\circ\psi$, we obtain
[…]
However, the vector field $\psi_{*}v\perp v$ is homotopic to $v$ through nowhere vanishing vector fields on $\mathrm{S}^{1}\times\mathrm{S}^{1}$ and therefore $\sigma(j,\psi_{*}v)=\sigma(j,v)$ by (2.6) and (2.5). Hence, $\sigma(j,v)$ is unlinked with $j(\ell)$ too. ∎
Corollary 2.9.
If $\sigma(j,v)\neq 0$ for a Lagrangian embedding $j:\Sigma_{n}\to\mathbb{C}^{n}$, then $n$ is odd and $\sigma(j\circ\rho,v)=0$ for the diffeomorphism $\rho$ given by (2.4).
Proof.
$\rho_{*}v=-v$ and the claim follows from (2.7) and Theorem 2.7. ∎
3. Classification of embeddings
Let us start with the special case $n=3$. The isotopy classes of smooth embeddings $j:\Sigma_{3}\to\mathbb{R}^{6}\cong\mathbb{C}^{3}$ are indexed by the following two invariants [20, p. 648]. First, the Whitney invariant
$W(j)\in\mathrm{H}_{1}(\Sigma_{3};\mathbb{Z})\cong\mathbb{Z},$
which is essentially the difference class between the Haefliger–Hirsch maps (3.2) of $j$ and of the standard embedding in $\mathbb{R}^{2}\times\mathbb{R}^{3}\cong\mathbb{R}^{5}\subset\mathbb{R}^{6}$. Second, the Kreck invariant
$\eta(j)\in\mathbb{Z}/|W(j)|\mathbb{Z},$
where we have used the above identification $\mathrm{H}_{1}(\Sigma_{3};\mathbb{Z})\cong\mathbb{Z}$. The Kreck invariant generalises the Haefliger invariant [12] classifying smoothly knotted $3$-spheres in $\mathbb{R}^{6}$.
In particular and most importantly for us, the Kreck invariant of an embedding with $W(j)=\pm 1\in\mathbb{Z}$ takes values in $\{0\}$ and therefore all such embeddings with the same Whitney invariant are isotopic. (This case is singled out in Corollary (a) on p. 649 of [20].)
The Boéchat–Haefliger Invariant Lemma on p. 661 of [20] asserts that if $\xi$ is an unlinked section of the normal bundle of $j$, then
$2W(j)=\pm\mathop{\mathrm{PD}}\,e(\xi^{\perp})\in\mathrm{H}_{1}(\Sigma_{3};\mathbb{Z}),$ (3.1)
that is, $2W(j)$ is Poincaré dual to the Euler class of the orthogonal complement of $\xi$ in the normal bundle of $j(\Sigma_{3})$. The sign $\pm$ in (3.1) is fixed by orientation conventions as stated right before the Difference Lemma on p. 658 of [20].
For a Lagrangian embedding $j$, an unlinked normal bundle section is given by $\xi:=i\cdot dj(v)$ after possibly precomposing $j$ with the diffeomorphism (2.4) by Corollary 2.9. Multiplication by $i$ defines an isomorphism $\xi^{\perp}\cong v^{\perp}$, where the latter complement is taken in $T\Sigma_{3}$. The orthogonal complement to $v$ is just the subbundle tangent to the $\mathrm{S}^{2}$-fibres in $\Sigma_{3}$, so its Euler class equals twice a generator of $\mathrm{H}^{2}(\Sigma_{3};\mathbb{Z})$. Hence,
$W(j)=\pm 1\in\mathbb{Z}\cong\mathrm{H}_{1}(\Sigma_{3};\mathbb{Z})$
by formula (3.1), where the sign is again fixed by the orientation conventions and the identification $\mathrm{H}_{1}(\Sigma_{3};\mathbb{Z})\cong\mathbb{Z}$. Now the above remark about the Kreck invariant implies the following result for $n=3$.
Proposition 3.1.
All Lagrangian embeddings $\mathrm{S}^{1}\times\mathrm{S}^{2}\hookrightarrow\mathbb{C}^{3}$ are smoothly isotopic up to precomposition with (2.4).
Let now $n\geq 2$ be arbitrary. Trivialise the tangent bundle of $\Sigma_{n}$ using the dilation invariant vector fields
$\zeta_{k}=\|x\|\frac{\partial}{\partial x_{k}},\quad k=1,\ldots,n,$
on $\mathbb{R}^{n}\setminus\{0\}$ and note that
$v=\sum\frac{x_{k}}{\|x\|}\zeta_{k}.$
Fix also a complex linear trivialisation of $T\mathbb{C}^{n}$ so that $T_{a}\mathbb{C}^{n}\cong\mathbb{C}^{n}$ for all $a\in\mathbb{C}^{n}$.
Let $j:\Sigma_{n}\to\mathbb{C}^{n}$ be an embedding and let $\Sigma_{n}^{\circ}$ denote the complement of a small $n$-ball in $\Sigma_{n}$. There exists a unique up to homotopy normal vector field $\nu_{j}$ on $j(\Sigma_{n})$ such that it does not vanish on $\Sigma_{n}^{\circ}$ and the pushoff of $j(\Sigma_{n})$ by $\nu_{j}$ is nullhomologous in $\mathbb{C}^{n}\setminus j(\Sigma_{n}^{\circ})$, see [13, p. 134]. (This vector field is called the associated vector field of $j$ in [3, §5.1] and the Haefliger–Hirsch vector field in [6, §6.2]. It is also the unlinked section of the normal bundle from [20] appearing in (3.1).)
Specialising [13, Theorem (2.2)] and the remarks right after it to our situation, we see that for $n\geq 4$ the isotopy classes of embeddings $j$ are in one-to-one correspondence with the homotopy classes of their Haefliger–Hirsch maps
$\Sigma_{n}^{\circ}\ni x\longmapsto\bigl{(}dj_{x}(\zeta_{1}(x)),\dots,dj_{x}(\zeta_{n}(x)),\nu_{j}(j(x)\bigr{)}\in\mathcal{V}_{2n,n+1}$ (3.2)
to the Stiefel manifold of real $(n+1)$-frames in $\mathbb{R}^{2n}\cong\mathbb{C}^{n}$, cf. [3, §5.2]. (It is convenient to use arbitrary frames here but it is well-known that $\mathcal{V}_{*,*}$ retract onto Stiefel manifolds of orthonormal frames [21, §2].)
Since $\mathcal{V}_{2n,n+1}$ is $(n-2)$-connected [21, Satz 8] and $\Sigma_{n}^{\circ}$ retracts onto the wedge sum $\mathrm{S}^{1}\vee\mathrm{S}^{n-1}$, it follows that for $n\geq 3$ the homotopy class of (3.2) is determined by the image of the meridional $(n-1)$-sphere $m\subset\Sigma_{n}^{\circ}$ in $\pi_{n-1}(\mathcal{V}_{2n,n+1})$. The latter group is isomorphic to $\mathbb{Z}$ if $n$ is odd and to $\mathbb{Z}/2\mathbb{Z}$ if $n$ is even [21, Satz 9], which completes the classification of smooth embeddings of $\Sigma_{n}$ in $\mathbb{C}^{n}$ for $n\geq 4$.
Assume now that $j$ is a Lagrangian embedding. By Corollary 2.9, up to precomposing $j$ with the reflection (2.4) for odd $n$, we may choose $\nu_{j}=i\cdot dj(v)$ so that the map (3.2) takes the form
$x\longmapsto\left(dj_{x}(\zeta_{1}(x)),\dots,dj_{x}(\zeta_{n}(x)),i\sum\limits_{k=1}^{n}\frac{x_{k}}{\|x\|}dj_{x}(\zeta_{k}(x))\right).$
Furthermore, since a Lagrangian embedding is totally real, the vectors $dj_{x}(\zeta_{k}(x))$ form a complex frame of $\mathbb{C}^{n}$ for every $x\in\Sigma_{n}^{\circ}$. Therefore there is a smooth map
$\Phi:\Sigma_{n}^{\circ}\longrightarrow\mathrm{GL}(n,\mathbb{C})$
such that
$dj_{x}(\zeta_{k}(x))=\Phi(x)e_{k},$
where $e_{k}$, $k=1,\ldots,n$, is our fixed complex frame of $\mathbb{C}^{n}$. Hence, the map (3.2) can be written as
$x\longmapsto\left(\Phi(x)e_{1},\dots,\Phi(x)e_{n},\Phi(x)\sum\limits_{k=1}^{n}\frac{x_{k}}{\|x\|}ie_{k}\right).$ (3.3)
The action of $\mathrm{GL}(n,\mathbb{C})\subset\mathrm{GL}_{+}(2n,\mathbb{R})$ on $\mathcal{V}_{2n,n+1}$ induces a trivial map on the relevant homotopy group $\pi_{n-1}$ for all $n\neq 2,4$ by the proof of [6, Lemma 6.14]. Specifically,
  • (1)
    if $n$ is odd, then $\pi_{n-1}(\mathrm{GL}(n,\mathbb{C}))=\pi_{n-1}(\mathrm{U}(n))=0$;
  • (2)
    if $n\neq 2,4,8$ is even, then the map
    $\pi_{n-1}(\mathrm{GL}_{+}(2n,\mathbb{R}))\longrightarrow\pi_{n-1}(\mathcal{V}_{2n,n+1})=\mathbb{Z}/2\mathbb{Z}$
    induced by the action is trivial;
  • (3)
    if $n=8$, then the map in (2) is surjective but the map
    $\mathbb{Z}=\pi_{7}(\mathrm{GL}(8,\mathbb{C}))\longrightarrow\pi_{7}(\mathrm{GL}_{+}(16,\mathbb{R}))=\mathbb{Z}$
    is the multiplication by $2$, as observed already in [3, §5.3].
Thus, for $n\neq 2,4$, the Haefliger–Hirsch map (3.3) of any Lagrangian embedding (precomposed with (2.4) if $n$ is odd and $\sigma(j,v)\neq 0$) is homotopic to the standard map
$x\longmapsto\left(e_{1},\dots,e_{n},\sum\limits_{k=1}^{n}\frac{x_{k}}{\|x\|}ie_{k}\right),$ (3.4)
which proves the following result for all $n\geq 5$.
Proposition 3.2.
All Lagrangian embeddings $\mathrm{S}^{1}\times\mathrm{S}^{n-1}\hookrightarrow\mathbb{C}^{n}$, $n\geq 5$, are smoothly isotopic up to precomposition with (2.4) if $n$ is odd.
Remark 3.3.
If $n=4$, the following argument from [3, p. 232] shows that both isotopy classes of embeddings $j:\Sigma_{4}\to\mathbb{C}^{4}$ have Lagrangian representatives. Note first that (for any $n\geq 3$) the action of the map $\Phi$ in (3.3) on $\pi_{n-1}(\Sigma_{n}^{\circ})$ can be chosen at will by [1, §6.2] and the $h$-principle for Lagrangian immersions of $\mathrm{S}^{n-1}$ [4, §24.3]. However, the map
$\mathbb{Z}=\pi_{3}(\mathrm{GL}(4,\mathbb{C}))\longrightarrow\pi_{3}(\mathcal{V}_{8,5})=\mathbb{Z}/2\mathbb{Z}$
is surjective and so each homotopy class of maps $\Sigma_{4}^{\circ}\to\mathcal{V}_{8,5}$ can be represented by the Haefliger–Hirsch map of a Lagrangian embedding.
Remark 3.4.
For $n=3$, the proof of Proposition 3.2 amounts to an alternative computation of the Whitney invariant. Let $\mathcal{N}\to\mathcal{V}_{2n,n+1}$ be the tautological normal bundle. (The fibre of $\mathcal{N}$ at a frame is the orthogonal complement to the span of the frame.) An orientation on $\mathbb{R}^{2n}$ defines an orientation on $\mathcal{N}$. Two observations are immediate:
  1. (1)
    The pull-back of $\mathcal{N}$ by the Haefliger–Hirsch map (3.2) is the orthogonal complement $\nu_{j}^{\perp}(=\xi^{\perp})$ in the normal bundle of $j$.
  2. (2)
    The pull-back of $\mathcal{N}$ to the meridional $2$-sphere $m=\{\|x\|=1\}$ by the standard map (3.4) is the tangent bundle of $m$.
For a Lagrangian embedding $j$ (precomposed with (2.4) if needed), the maps (3.2) and (3.4) are homotopic. Hence, $\langle e(\xi^{\perp}),[m]\rangle=2$ and therefore $W(j)=\pm 1\in\mathbb{Z}\cong\mathrm{H}_{1}(\Sigma_{3};\mathbb{Z})$ by (3.1).
It remains to show that precomposition with the involution $\rho$ of $\Sigma_{n}$ defined by (2.4) does indeed change the smooth isotopy class of a Lagrangian embedding for odd $n$. Swapping $j$ and $j\circ\rho$ if necessary, we may assume that $\sigma(j,v)=0$ by Corollary 2.9 so that $\sigma(j\circ\rho,-v)=0$ by (2.7). Hence, the argument leading to Proposition 3.2 shows that the Haefliger–Hirsch maps of $j$ and $j\circ\rho$ are homotopic to
$x\longmapsto\left(e_{1},\dots,e_{n},\pm\sum\limits_{k=1}^{n}\frac{x_{k}}{\|x\|}ie_{k}\right)$ (3.5)
with the $+$ sign for $j$ and the -- sign for $j\circ\rho$. The restrictions of the maps (3.5) to the meridional sphere $m=\{\|x\|=1\}$ parametrise the (homotopy) fibre $\mathrm{S}^{n-1}\sim\mathcal{V}_{n,1}\times\mathbb{R}^{n}$ of the ‘forgetful’ fibre bundle
$\mathcal{V}_{2n,n+1}\longrightarrow\mathcal{V}_{2n,n}$
and differ by the antipodal map of $\mathrm{S}^{n-1}$. For odd $n$, the antipodal map is orientation reversing and hence acts non-trivially on $\pi_{n-1}(\mathrm{S}^{n-1})=\mathbb{Z}$. Furthermore, $\pi_{n-1}(\mathcal{V}_{2n,n})=0$ by [21, Satz 8] and $\pi_{n}(\mathcal{V}_{2n,n})=\mathbb{Z}/2\mathbb{Z}$ for odd $n>1$ by [21, Satz 9], so the homotopy exact sequence
$\pi_{n}(\mathcal{V}_{2n,n})\to\pi_{n-1}(\mathrm{S}^{n-1})\to\pi_{n-1}(\mathcal{V}_{2n,n+1})\to\pi_{n-1}(\mathcal{V}_{2n,n})$
shows that the embedding of the fibre is an isomorphism on $\pi_{n-1}$. Hence, the two maps in (3.5) are not homotopic, which proves the next statement.
Proposition 3.5.
A Lagrangian embedding $\mathrm{S}^{1}\times\mathrm{S}^{n-1}\hookrightarrow\mathbb{C}^{n}$ with odd $n\geq 3$ is not smoothly isotopic to its composition with (2.4).
Theorem 1.1 follows by combining Propositions 3.1, 3.2, and 3.5.
Acknowledgements
The author is grateful to Leonid Polterovich, Felix Schlenk, and Arkadiy Skopenkov for their comments on the first draft of this paper.